40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 17137     Entry time: Thu Sep 8 16:03:25 2022
Author: Yehonathan 
Type: Update 
Category: LSC 
Subject: Realignment, arm locking, gains adjustments 

{Anchal, Yehonathan}

We came this morning and the IMC was misaligned. The IMC was realigned and locked. This of course changed the input beam and sent us down to a long alignment journey.

We first use TTs to find beam on BHD DCPD/Camera since it is only single bounce on all optics.

Then, PR2/3 were used to find POP beam while keeping the BHD beam.

Unfortunately, that was not enough. TTs and PRs have some degeneracy which caused us to lose the REFL beam.

Realizing this we went to AS table to find the REFL beam. We found a ghost beam that decieved us for a while. Realizing it was a ghost beam, we moved TT2 in pitch, while keeping the POP DCPD high with PRs, until we saw a new beam on the viewing card.

We kept aligning TT1/2, PR2/3 to maximize the REFL DCPD while keeping the POP DCPD high. We tried to look at the REFL camera but soon realized that the REFL beam is too weak for the camera to see.

At that point we already had some flashing in the arms (we centered the OpLevs in the beginning).

Arms were aligned and locked. We had some issue with the X-ARM not catching lock. We increased the gain and it locked almost immediately. To fix the arms gains correctly we took OLTFs (Attachment) and adjusted the XARM gain to 0.027 to make the UGF at 200Hz.


Both arms locked with 200 Hz UGF from:

From GPS: 1346713049
    To GPS: 1346713300

From GPS: 1346713380
    To GPS: 1346714166

HEPA turned off:
From GPS: 1346714298
    To GPS: 1346714716

 

ELOG V3.1.3-